# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 3 Supervised Learning-based CMR
Existing supervised learning-based methods for CMR can be categorized into three key dimensions: data, model architecture, and optimization objectives. First, as discussed in Section 2, the construction and annotation of multimodal triplet data pose significant challenges. Consequently, research in this area primarily focuses on automating data augmentation and addressing annotation noise. The goal is to develop techniques that can efficiently generate large-scale triplet datasets while ensuring the quality and reliability of annotations. Second, in terms of model architecture, existing studies investigate various strategies for integrating visual and linguistic modalities. These efforts emphasize the precise modeling and fusion of semantic information, focusing on retaining, modifying, and supplementing semantics during integration. The goal is to design models that effectively align and fuse both modalities, ensuring that the combined representation captures the intended modifications and preserves the relevant context. Finally, the optimization objectives under the triplet data framework represent another critical aspect. Building on contrastive learning, recent advancements have introduced improvements to objective functions, aiming to enhance the alignment between the fused input and the visual target, thereby improving retrieval performance. In the following, we provide a detailed review of the progress made in these areas, highlighting key contributions and ongoing challenges.
3.1 Data Augmentation
@@
The quantity and quality of data play a crucial role in deep learning, especially in light of the empirical evidence supporting scaling laws. Unlike conventional cross-modal visual-language retrieval, where image-text pairs or video-text pairs can be readily collected from the web, CMR requires training data in the form of triplets, e.g., {reference image, relative text, target image}. In such triplets, the reference image and relative text jointly represent the user’s retrieval intent, with the relative text specifying flexible modification requests grounded in the rich visual objects and attributes of the reference image. The target image serves as the desired output of the retrieval process, with the relative text effectively describing the semantic differences between the reference and target images. However, in practice, triplets that capture such nuanced semantic differences are scarce, making constructing large-scale, high-quality triplet data a fundamental challenge in this field.
Early works primarily relied on manual annotation to construct triplet datasets. While these efforts provided a foundation for dataset construction, the manual annotation process is time-consuming and labor-intensive, leading to several significant limitations: (1) Limited Data Scale: The size of datasets is constrained by the human effort required for annotation; (2) Restricted Data Domains: Datasets are often confined to specific domains, limiting their generalizability; (3) Simplistic Relative Text Descriptions: The modifications expressed in the relative text are frequently simplistic, lacking the complexity required for real-world applications; (4) Annotation Noise and Uncertainty: Human annotations may introduce noise, which becomes a bottleneck for further advancements in the field.
To mitigate these issues, current work focuses on leveraging image and text generation techniques to automate the construction of triplet data. These approaches aim to improve the quantity of triplet data while having high-quality triplet annotations by minimizing the introduction of noise and uncertainty. Below, we highlight related key works and their technical innovations.
3.1.1 Automatic Data Construction
Works in this research line can be divided into two major categories according to the type of generation modality: (1) generating language content only  [1, 2, 3, 4, 5, 6, 7]; and (2) generating language and visual content simultaneously [8, 9, 10].
Generating Language Content: The augmentation of triplet data through text generation can be categorized into two main approaches: 1) generating text from a single image and 2) generating text describing differences between two images. These methods generate relative descriptions from a single image, typically by modifying its caption or masking parts of the input information. Liu et al. [1] develop a pipeline using large-scale image-text datasets (e.g., Laion-COCO) to generate relative descriptions by editing original captions. This is achieved using predefined templates (e.g., quantity changes, comparisons) or large language models (LLMs) like ChatGPT. The generated text is compared with candidate image descriptions to select the most similar target image, forming triplets. Two datasets are constructed: Laion-CIR-Template (template-based) and Laion-CIR-LLM (LLM-based), each containing approximately 16K triplets. The authors note that template-generated relative descriptions tend to be more effective, especially in fine-grained retrieval tasks, compared to those generated by LLMs. Hou et al. [2] propose a masked training strategy, which involves randomly masking patches of an image to create a reference image, while the original image serves as the target. An image caption generator is then used to produce a caption for the original image, which acts as the modification text, creating pseudo triplets. Zhang et al. [3] focus on generating challenging negative examples. They modified query sentences or replaced reference images with semantically mismatched alternatives to create pseudo triplets with subtle differences. This improved the model’s ability to distinguish examples and learn a better metric space for complex retrieval tasks. Liu et al. [4] introduce bidirectional training, where the model is trained on both forward (reference image + modification text to target image) and backward (target image + reverse modification text to reference image) queries. This expands data volume and strengthens multimodal fusion capabilities.
These methods generate descriptions of differences between two images to form triplets. Jiang et al. [5] propose a synthetic label generation strategy that consists of three steps: (1) selected visually similar image pairs, (2) generated difference descriptions for the pairs using pre-trained visual-language models (e.g., BLIP) and LLMs, and (3) filtered the generated text for semantic quality. This ensured high-quality triplet construction. Jang et al. [6] propose the Visual Delta Generator (VDG), which first searches for reference images and their corresponding target images in auxiliary datasets, ensuring that the image pairs have both similarities and differences. Then, the VDG generates text descriptions of the visual differences between the reference and target images, constructing pseudo triplets. Levy et al. [7] introduce LaSCo, a large-scale dataset that is ten times the size of existing CMR datasets. They selected visually similar image pairs with differing answers to the same question and used GPT-3 to generate transition text describing the changes. This approach resulted in a large-scale dataset LaSCo, significantly expanding available triplet data and boosting model performance.
Generating Language and Visual Content: Some works not only generate text but also generate images to further augment the triple data. Zhang et al. [8] propose CTI-IR, a GAN-based network that jointly trains generative and retrieval models to handle image generation and retrieval tasks. The model uses global-local collaborative discriminators to ensure semantic consistency between the generated image and the modification text by capturing both overall and fine-grained differences. Gu et al. [9] introduce a diffusion-based model for constructing a large-scale synthetic dataset called SynthTriplets18M. By leveraging existing image-caption datasets (e.g., COYO 700M, LAION-2B-en-aesthetic), they first generate diverse modified texts by replacing object terms in the original captions or editting operation via LLMs. This process creates a wide range of semantic variations, covering more visual concepts. Then, a latent diffusion model is used to generate high-quality target images based on the reference image and modification text. This approach enhances the model’s ability to handle diverse modifications and introduces capabilities like handling negative text and image mask conditions. Zhou et al. [10] propose one strategy, generating multi-modal triplets from large-scale image-text datasets. They use LLMs (GPT 3.5) to generate modification texts, which can be to change certain objects, colors, positions, etc, and then employ stable diffusion to create target images, ensuring alignment between the modified text and the generated image.
3.1.2 Noise/Uncertainty in Data
In addition to the volume of data, ensuring data quality is a critical focus in existing research. High-quality triplet data can significantly enhance model training effectiveness by mitigating the impact of noisy data, thereby improving retrieval performance. Data uncertainty and noise primarily manifest in two forms: (1) Content noise, where erroneous content is incorrectly deemed semantically relevant [5, 1, 9, 11, 2, 7]. For instance, the target image does not fully align with the retrieval intent specified by the modification text. This issue is particularly pronounced in automatically generated data. (2) Annotation uncertainty, where the target image in the triplet is not the sole positive example, as other images in the dataset may also satisfy the retrieval intent defined by the reference image and modification text [12, 13, 14, 15, 16, 17]. Existing research has addressed these challenges through a series of studies. Below, we review and summarize these efforts in detail.
Content Noise: Existing methods primarily employ threshold-based filtering strategies to ensure semantic relevance and consistency. These strategies focus on measuring the similarity between reference images, modified texts, and target images to align with the intended modification intention. Jiang et al. [5] and Liu et al. [1] both utilize semantic similarity filtering to ensure consistency between generated modification texts and image pairs. Jiang et al. [5] screen modification texts within the language space, retaining triplets with high semantic similarity to both reference and target images. Similarly, Liu et al. [1] generate modification instructions using a large language model, followed by sentence transformation to produce edited descriptions. They apply a similarity-based filter to select triplets with high semantic relevance between edited descriptions and reference images. Gu et al. [9] design a CLIP-based filtering strategy, calculating CLIP similarity between reference and generated target images, as well as between original and modified texts, and applying thresholds to ensure visual and semantic consistency. They introduce directional CLIP similarity to evaluate alignment between image and text modifications and verify that generated images reflect specified keywords or instructions.
Lin et al. [11] introduce a similarity-based data augmentation method. Specifically, they utilize CLIP’s image encoder to extract features from all images in the training set, and then compute the cosine similarity between the reference or target images and other images in the training set. Based on the similarity scores, several of the most similar images are selected as new reference or target images, thereby generating pseudo-triplets. To ensure quality, the generated pseudo-triplets select the most similar images, constrained by predefined similarity thresholds and a limit on the number of pseudo-triplets per image to control quality and prevent overfitting. Hou et al. [2] introduce a random sampling strategy based on the 3-$\sigma$ rule to select challenging samples for fine-tuning. They generate pseudo-modification texts for unlabeled image pairs, calculate distance metrics, and define a candidate range using the 3-$\sigma$ rule. Randomly selected samples from this range mitigate noise while ensuring sufficient challenge, enhancing model robustness. Levy et al. [7] develop an analytical tool to detect redundant noise in queries by evaluating the contribution of different modalities. Queries with minimal impact from either image or text are identified as containing redundant information. By identifying and filtering out triplets that include such redundant information, the quality of the generated data can be enhanced.
Annotation Uncertainty: To address the issue of annotation uncertainty in triplets, where a single query may correspond to multiple valid target images, existing methods focus on mitigating false positives and enhancing semantic diversity. These approaches can be categorized into two main strategies: uncertainty-aware modeling and constructing evaluation datasets with multiple ground-truth labels.
Some works focus on modeling uncertainty in data. Wen et al. [12] propose a target similarity distribution-based matching regularization module to handle false positives. This module calculates the similarity between the ground-truth target image and other candidates within a mini-batch, dynamically adjusting match scores to allow flexible handling of partially matching candidates rather than treating all non-target images as negatives. Zhang et al. [13] address triplet uncertainty by introducing the Consensus Network (Css-Net), which employs four distinct compositors to generate diverse image-text embeddings. This reduces reliance on any single potentially biased compositor and promotes consistent outputs through Kullback-Leibler divergence loss. In evaluation, decisions from the compositors are combined via a weighted scheme to enhance consistency. Chen et al. [14] propose a rank-aware uncertainty method to model many-to-many correspondences. Their approach includes three modules: (1) In-sample Uncertainty, which represents reference and target image features as Gaussian distributions to generate multiple potential matching candidates; (2) Cross-sample Uncertainty, which identifies potential positive samples by comparing similarity distributions across different samples; and (3) Distribution Regularization, which minimizes discrepancies between source and target feature distributions to improve cross-modal alignment. Chen et al. [15] further enhance uncertainty modeling by introducing fluctuations in the feature space that simulate multi-granularity queries. The model uses one-to-one matching, while for coarse-grained queries, it employs one-to-many matching within an uncertainty range. An uncertainty regularization module dynamically adjusts matching targets based on query precision, reducing false positives and improving retrieval accuracy.
Some works propose new evaluation datasets that annotate multiple ground truth labels for each query. Baldrati et al. [16] introduce the CIRCO dataset, which addresses the limitation of single ground truth labels by annotating multiple valid target images for each query. Using multimodal retrieval models, the top 100 relevant images are retrieved for each query, filtered to 50 visually similar images, and annotated to include all valid matches. This results in an average of 4.53 ground truth labels per query, reducing false negatives and better simulating real-world complexity. Similarly, the CFQ (Challenging Fashion Queries) dataset [17] introduces multi-label annotations, where each query includes multiple positive and negative labels. Annotators select target images from the same category that match the query as positive labels and those that do not as negative labels. This provides a more comprehensive and challenging benchmark for evaluating text-guided fashion image retrieval models.
@@ 3.2 Model Architecture
For input reference images and modification texts, the core challenge of combined multimodal retrieval lies in how to integrate the two modalities, that is, to represent a composite semantics that is similar to the reference image but adjusted in detail according to the modification text. Existing works can be categorized into three main aspects: (1) The most common approach is to design feature combiners for visual and language modalities, fusing the two modalities to achieve modified multimodal semantic representation; (2) Plug-and-play components, which offer broad compatibility across different existing methods and hold significant summarizing value; (3) Non-combiner methods that take an alternative path, implementing combined multimodal retrieval through novel architectures, providing important inspiration.

### v2

@@ 3 Supervised Learning-based CMR
Existing SL-CMR can be categorized into three key dimensions: data, model architecture, and optimization objectives. First, as discussed in Section 2, the construction and annotation of multimodal triplet data pose a fundamental challenge in this field. Second, in terms of model architecture, existing studies investigate various strategies for combining visual and linguistic modalities. These efforts emphasize the precise semantic modeling, focusing on retaining, modifying, and supplementing semantics during combination. The goal is to ensure that the combined representation captures the intended modifications and preserves the relevant context. Finally, the optimization objectives in the triplet data framework represent another critical aspect. Building on contrastive learning, recent advancements have introduced improvements to objective functions. In the following, we provide a detailed review of the progress made in these areas, highlighting key contributions and ongoing challenges.
3.1 Data Construction
@@
The quantity and quality of data play a crucial role in deep learning, especially in light of the empirical evidence supporting scaling laws [152]. Unlike conventional cross-modal visual-language retrieval, where image-text pairs or video-text pairs can be readily collected from the web, CMR requires training data in the form of triplets, e.g., {reference image, modified text, target image}. In such triplets, the reference image and modified text jointly represent the user’s retrieval intent, with the text specifying flexible modification requests grounded in the rich visual objects and attributes of the reference image. The target image serves as the desired output of the retrieval process, with the relative text accurately describing the semantic differences between the reference and target images. However, in practice, triplets that capture such nuanced semantic differences are scarce, making construction of large-scale, high-quality triplet data a fundamental challenge in this field.
Early works primarily relied on manual annotation to construct triplet datasets. CSS [20] introduces a simple 3D scene rendering dataset. They first generate relative modification texts for reference images based on manually defined templates. These modifications target different object attributes such as color, shape, and size, and involve operations such as addition, removal, and alteration, for example, “add a red cube”. The CLEVR toolkit [153] is then used to synthesize new target images, thereby constructing triplet data with relative semantic differences. CIRR [21] is the first large-scale benchmark tailored for composed image retrieval with real-world photos. Derived from the NLVR2 dataset [22], CIRR includes 21,552 images and 36,554 triplets. The dataset emphasizes semantic diversity and minimizes false positives, with a single target image per query and an online evaluation protocol. CIRCO [23] introduces multiple ground truths per query (average of 4.53), which enhances annotation quality and reduces false negatives. Another earlier benchmark is Birds-to-Words [24], which contains 3,347 bird image pairs from iNaturalist, each accompanied by an average of 4.8 descriptive paragraphs detailing fine-grained differences. Though limited in size, it offers richer linguistic content per instance. MIT-States [25] comprises 60K images labeled with 249 object nouns and 115 state adjectives, supporting adjective change scenarios (such as “new camera”, “red tomato”), emphasizing the combination generalization ability.
In addition, representative benchmarks also include the fashion images domain. FashionIQ [26] serves as a representative benchmark for composed fashion image retrieval. It comprises 77,684 images across three categories, i.e., dresses, shirts, and tops & tees, including 60,272 triplets. In contrast, Fashion200k [27] contains over 200,000 product images sourced from online shopping platforms, with textual descriptions filtered and cleaned to retain 4,404 unique attributes for joint embedding. However, it lacks natural modification sentences, making it more suitable for attribute-based CMR. The Shoes dataset [28] consists of 10,751 triplets from like.com, annotated with fine-grained relative descriptions.
Benchmarks in other domains, including videos [29, 30, 31], remote sensing images [109, 110], person images [111], sketch images [154, 155, 156, 157, 158], multi-turn conversations [115, 117, 114], etc., can be referred to Sec. 6. Although the above datasets provide a feasible way to construct triplet data, the manual annotation process is time-consuming and laborious, which also leads to some key issues, such as the small data volume, single data domain, relatively simple text description, uncertainty, and noise in annotations. These problems restrict the further development of this field. Subsequent work has carried out a new paradigm (see Sec. 5) for automatic data generation to alleviate this problem.
@@ 3.2 Model Architecture
For input reference images and modification texts, the core challenge of composed multi-modal retrieval lies in how to combine the two modalities, i.e., to represent a composite semantics that is similar to the reference image but adjusted in detail according to the modification text. Existing works can be categorized into three main aspects: (1) The most common approach is to design feature combiners for visual and language modalities, fusing the two modalities to achieve modified multimodal semantic representation; (2) Plug-and-play components, which offer broad compatibility across different existing methods and hold significant generalization value; (3) Non-explicit combiner methods that take an alternative path, implementing composed multimodal retrieval through novel architectures, providing important inspiration.
Figure 3: The existing multimodal combiner methods for CMR can be classified into 6 categories, progressing from simple to sophisticated structures.